Solve (x + 5)(x – 3) = -2(2x + 7)
Answer by nyc_function(2741)

(x + 5)(x – 3) = -2(2x + 7)
I will start and you can finish.
x^2 + 2x - 15 = -4x - 14
x^2 + 2x + 4x - 15 + 14 = 0
x^2 + 6x - 1 = 0
Can you continue now?